OpenStudy (anonymous):

Determine the interval of convergence for the Taylor series about the indicated point for the following function [You do not need to find the Taylor series.]

OpenStudy (anonymous):

\[\frac{ 1 }{ x ^{3}+1}\] a=3

OpenStudy (anonymous):

this function is undefined at \(x=-1\) so the radius of convergence (i think) will be 4

OpenStudy (anonymous):

that's what i thought till i saw this\[x^3=cis 5\pi thus x=\frac{ 1 }{ 2 }+ \frac{ \sqrt{3}j }{ 2 } or -1 or \frac{ 1 }{ 2 }- \frac{ \sqrt{3}j }{ 2 } \]

OpenStudy (anonymous):

i could be wrong, but clearly this is undefined at \(x=-1\) so if you expand about 3 you cannot go more than 4 away

OpenStudy (anonymous):

not sure where it all came from

OpenStudy (anonymous):

those are the cubed roots of \(-1\)

OpenStudy (anonymous):

since this is a function of real variables, this has nothing to do with the problem those are the real and complex solutions to \(x^3+1=0\)

OpenStudy (anonymous):

so are you saying that the complex roots of the function should not be included?

OpenStudy (anonymous):

no they should not

OpenStudy (anonymous):

thx,

OpenStudy (anonymous):

yw